OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 3m - 5 = 7m - 21

OpenStudy (anonymous):

3m-7m=-21+5 -4m=-16 4m=16 m=4

OpenStudy (anonymous):

take 7m to the left side and -5 to the right side( remember the sign convention) 3m-7m = -21+5 -4m = -16 m = 4

OpenStudy (anonymous):

7m-21=3m-5 7m-3m=-5+21 4m=16 m= 16/4 =4